CALL FOR PAPERS — ICNNAI’2006

4th International Conference on Neural Networks and Artificial Intelligence

31 May - 2 June, 2006, Brest, Belarus

Official Language: English Paper Submission deadline: 8 March 2006

(http://neuro.bstu.by/icnnai2006/cfp.html)

SCOPE AND AIM:

A realization of Artificial Intelligence (AI) by Neural Networks (NN) might sound somehow old but will be an ever lasting research area.

This conference intends to provide an interactive forum for discussions, debates and brainstorms among researchers interested in combining these two fields, in a creative, exciting, and free atmosphere. In the process, we expect collaborations among those who have complementary ideas.

The conference will feature three plenary talks by - Heinz Muhlenbein ... Fraunhofer Institute (Germany);

- Robert Hiromoto ... University of Idaho (US);

- Colin Frayn ... University of Birmingham (UK). oral sessions with accepted papers and a couple of panel discussions are also planned.

We solicit papers of the latest techniques and achievements from theoretical, experimental, as well as engineering aspect in all areas of Neural Networks aiming Artificial Intelligence.

TOPICS:

The conference covers all topics in Artificial Intelligence by Neural Networks. Or far beyond, though we dare not to use a cliche like ”including but not limited to” such as

- Brain-Inspired Information Technology;

- Modeling Learning and Memory;

- Robotics and Sensory Systems and Motor Controls;

- Robotics and Cognition, Emotion and Behavior;

- Brain-like Knowledge Extraction or Recovery in Databases;

- Evolving brain-like structure; and

- Artificial Life. Or their fusion. Note again that the list is far from exhaustive nor exclusive, which are also cliche though.

SUBMISSIONS:

The papers must be in English with the maximum length of 8 pages, following the guidelines in the authors instructions which can be downloaded at http://neuro.bstu.by/icnnai2006/authkit.html where you can find a Style-file for LaTeX, as well as a skeleton for MS-WORD.

All submissions will be reviewed by the respective international program committees.

Papers will be selected based on their originality, timeliness, significance, relevance and clarity of presentation (forgive me a cliche again).

Please, however, do not expect reviewers to be a proof-reader. High quality submission is the author’s responsibility and inadequately prepared papers may be rejected.

Papers should be submitted in PDF format

(1) via the ONLINE SUBMISSION SYSTEM linked in the page to

http://neuro.bstu.by/icnnai2006/lite/prs.php

or

(2) by attaching it to an e-mail to icnnai@bstu.by with the subject ICNNAI SUBMISSION and with the attached file-name being ¡author-full-name¿.pdf.

PLEASE DON’T FORGET TO INCLUDE AT MOST 5 KEY-WORDS WHICH APPROPRIATELY REPRE-SENT WHAT TOPICS YOUR PAPER EXPLORES. (if you submit the paper by email these key-words should be in the text of e-mail , and if you submit it via the web-page simply fill in the box. This is for the purpose of efficiency in reviewing process.)

Receipt of submitted paper will be notified within a couple of working days after the deadline.

After review process of all submissions, you will be notified by 29 March whether your contribution is accepted or not.

Please notice that at least one of the authors of an accepted paper must register for the conference and present the paper.

The final program of the conference will be appeared on the conference web-page

http://neuro.bstu.by/icnnai2006/cfp.html

with the other up-to-date information.

KEY DATES:

The time schedule for the publication of the proceedings is as follows:

- Submission deadline: 08 March 2006

- Notification of acceptance: 29 March 2006

- Camera ready copy due: 19 April 2006

Please notice that we have prolonged these schedule from the ones in the preliminary Call for Paper. Please do not expect the schedule to be extended any more. Reviewers will have three weeks to review: accepted author will have three weeks to improve the paper, and we have six weeks to print the proceedings. Each of these three due dates will be a MUST.

PRESENTATION:

To keep the workshop lively we intend to keep the time of individual presentations efficiently. We plan a 20 minutes of oral presentation for all accepted papers.

Accepted papers will be published in the proceedings which will be available at the conference.

REGISTRATION FEE:

Registration fee is 125 USD for early registration which is to be done before by 19 April, after the date, including registration on site, the fee will be 175 USD.

The fee includes:

- proceedings;

- one welcome cocktail party (planned on 30 May in the evening);

- one reception party (planned on 31 May in the evening); and

- snacks, soft drinks at coffee-breaks between the sessions.

A limited number of the support to students from CIS countries will be also considered. Feel free to contact us.

We now are preparing for the bank-account for early registration. Please check it out later in up-to-date information at the conference web page.

VENUE:

The city of Brest is in a stone throw away from the border to Poland, and situated 200km from Warsaw and 300km from Minsk. From either of these cities we can access to the city by train quite easily.

Brest is a wonderful place to live, along with historical places which are outstanding for a city with medium size. It is a terrific place, personally speaking though, if you are interested in great nature.

We have the Brest fortress which is the main attraction of the city. The fortress made a prolonged defence against Germans in 1941 and was ruined in the result of severe battles. After the war the remains of the fortress were turned into a grandiose memorial to its defenders. (From the web-site bellow. Sorry for that but without a permission though.)

We have what they call largest forest in Europe in which we might visit the gorgeous Summer-House once presidents of former Soviet Union stayed during the summer holiday.

We have other Historical places as well as fancy Museums.

If you find time to walk to a big open market which is located in a walking distance from the conference site, you will enjoy various kinds of real fresh vegetables and fruits grown in Brest region. Life in Brest is wonderful. We hope to see you in Brest.

More about the city Brest is at

http://brestonline.com/en/info/index.html,

Also enjoy

http://www.reference.com/browse/wiki/Brest

HOTEL INFORMATION:

We have a couple of good Hotels. See, for example, http://www.belarustravel.by/en/Brest.

We also have a limited number of nice facilities for lodging in the campus of the Brest State Technical University where the conference will be held.

Later the more detailed list for the hotel accommodation will be appeared in the conference web-page.

Looking forward to receiving your paper(s) and meeting you at the conference in Brest.

HONORARY CHAIR:

• Petr Poyta (Belarus)

CONFERENCE CHAIR:

• Vladimir Golovko (Belarus)

CONFERENCE CO-CHAIRS:

• Rauf Sadykhov (Belarus)

• Akira Imada (Belarus)

INTERNATIONAL PROGRAM COMMITTEE:

• Juan Miguel Santos (Argentina)

• Helmut Mayer (Austria)

• Alexander Doudkin (Belarus)

• Vladimir Golenkov (Belarus)

• Vladimir Ptichkin (Belarus)

• Vladimir Rubanov (Belarus)

• Erinija Pranckeviciene (Canada)

• Andres Perez-Uribe (Chile)

• Xu Lisheng (China)

• Kurosh Madani (France)

• Jean-Jacques Mariage (France)

• Hubert Roth (Germany)

• Pasqual Daponte (Italy)

• Vincenzo Piuri (Italy)

• Hitoshi Hemmi (JAPAN)

• Qiangfu Zhao (Japan)

• Sung-Bae Cho (Korea)

• Irina Bausova (Latvia)

• Saulius Maskeliunas (Lithuania)

• Khalid Saeed (Poland)

• Alexander Galushkin (Russia)

• Vladimir Redko (Russia)

• Lipo Wang (Singapore)

• Emin German (Turkey)

• Bora Kumova (Turkey)

• Mykola Dyvak (Ukraine)

• Anatoly Sachenko (Ukraine)

• Voloddymyr Turchenko (Ukraine)

• Bogdan Gabrys (UK)

• Robert Hiromoto (USA)

LOCAL COMMITTEE:

• Yury Savitsky (Belarus)

• Sergei Besobrasov (Belarus)

• Svetlana Besobrasova (Belarus)

• Andrew Dunets (Belarus)

• Larisa Gorbashko (Belarus)

• Pavel Kochurko (Belarus)

• Yury Kozurko (Belarus)

• Leonid Machnist (Belarus)

• Nikolay Maniakov (Belarus)

• Piotr Selezniov (Belarus)

CONTACT ADDRESS:

Conference e-mail: iccnai@bstu.by

Conference Chair: Prof. Vladimir Golovko Co-chair: Prof. Akira Imada

Intelligent Information Technologies Department Brest State Technical University Moskowskaja 267, Brest, Belarus phone: +375-162-42-6321 fax: +375-162-42-2127